# Bridging Gaps, Building Futures: Advancing Software Developer Diversity and Inclusion Through Future-Oriented Research

### v1

@@ 4. Results: Avoiding Dystopia
In Section 3, we have outlined four scenarios on how SE industrial practice and higher education could look like in 2030. These scenarios do not have a direct connection to SE research. However, SE research can and should contribute to industrial practice and higher education in order to avoid the dystopian scenarios and instead achieve a future closer to the utopia described. To do so, we outline four orthogonal areas in which we believe SE research needs to evolve in the coming years.
@@
In Scenario S1, several SDDI metrics guided the company’s interventions toward a safer place for Jamie to develop their career. Similarly, Kris used various indicators in their educational setting in S2. If the university in Scenario S4 had monitored their students’ experience and well-being using suitable metrics, they would have noticed the lack of belonging and feeling of inclusion among students such as Moss. These three scenarios motivate the use of diversity metrics that can be used by practitioners and educators to better understand diversity in their environment and make decisions based on them. SE research can contribute to this area by developing appropriate Methodologies and Metrics.
@@
In Scenario S3, Ash’s experiences conflated frustrations due to multiple diversity dimensions: socioeconomic background, language background, and gender. In contrast, various intersectional backgrounds were effectively supported in Scenario S2 through personalized learning, which has been shown to reduce educational inequality and foster inclusion for learners from diverse cultures (Strekalova-Hughes et al., 2021). These scenarios highlight the importance of developing a better understanding of Intersectionality in SE research, that is, when multiple diversity aspects overlap.
@@
Both Scenarios S1 and S3 demonstrate how important it is that knowledge about SDDI be transferred to and from the SE field. For example, in Scenario S3, a lack of transfer of new ideas from outside means that traditional gender roles play a key role at the company, thus negatively impacting Ash’s sense of belonging. In Scenario S1, Jamie and their employer thrive due to the adoption of diversity initiatives. As such, there needs to be an increased push towards Knowledge Transfer between SE research, industry, education, government bodies, and societal groups.
@@
Finally, industrial practice and higher education do not operate in isolation, but are embedded in the larger socioeconomic context. For example, in Scenarios S1 and S3, national circumstances had a decisive effect on the workplace situation—in opposite directions. Therefore, it is imperative to use better existing Large-scale Socioeconomic Understandings developed by other disciplines in SE research to provide more inclusive environments in the workplace and in higher education.
@@ 4.1. Methodologies and Metrics
Methodologies and metrics are essential for scientific exploration and discovery (Creswell and Creswell, 2017). Methodologies entail systematic research approaches for acquiring new knowledge through various research methods for data collection and analysis (Wohlin and Runeson, 2021). Metrics are used to quantify a characteristic of a real-world entity (Basili, 1988). SE research, in particular, uses a variety of methods and metrics to provide empirical evidence to support software development processes and tools (Easterbrook et al., 2008). For example, previous work has used different techniques and measures to explore concepts related to SDDI, including surveys to understand developers’ sense of belonging (Trinkenreich et al., 2023, 2024), interviews to understand barriers for women in online programming communities (Ford et al., 2016), and studies mining GitHub repository to explore toxicity in open source software (Miller et al., 2022). However, while all these examples are related to SDDI, an underlying theory describing how they connect is missing. For example, how do we investigate toxicity as a barrier to minorities’ participation and belonging? “How” includes methods and metrics. Table 1 summarises potential benefits, harms and research directions for this theme.
@@
The focus of this theme is not only on what has been published in the field of SE around diversity and inclusion and with which methods—but also a look at multiple domains, that is, SE, educational psychology, and management, to investigate more how insights from these fields could be incorporated into the methodologies and metrics used to research SDDI in SE. Additionally, investing in mixed-methods research (Creswell and Creswell, 2017), utilizing qualitative and quantitative data more efficiently, is an important future goal. Weighting, timing, and mixing are aspects that must be considered before planning mixed-methods SDDI research.
@@
Weighting refers to the priority given to qualitative and quantitative data in research (Creswell and Creswell, 2017). The priority can be the same or favor one over another, depending on the researcher’s goals and the audience. Timing is about deciding whether to collect data sequentially in phases or concurrently (Creswell and Creswell, 2017). In sequential data collection, either qualitative or quantitative data collection can occur first, depending on the research goal. Concurrent data collection involves simultaneously gathering data for qualitative and quantitative analyses, which can be interesting in time-sensitive projects where contacting participants multiple times for data collection is not feasible (e.g., single surveys with open and closed-ended questions) (Trinkenreich et al., 2022a). Previous SDDI mixed-methods research used both concurrent and sequential mixed-methods research. Trinkenreich et al. (2022a) followed a concurrent mixed-methods research collecting data through a single survey to qualitatively uncover the challenges faced by women in software development teams and segment those challenges across demographics of age, caregiving responsibilities, marital status, and tenure. Examples of SDDI sequential mixed-methods research included surveys and mining software repositories studies in varying orders. Vasilescu et al. (2015) started with a survey on perceptions of team diversity and then mined a software repository to measure how team productivity and turnover are impacted by gender and tenure diversity. Following the opposite order, Prana et al. (2021) started mining software repositories to investigate differences in gender inclusion in projects across geographical regions, followed by a survey aimed at developers from the various regions about factors that can potentially contribute to differences in developer participation based on gender and geography worldwide.
@@
Mixing involves choosing how to integrate or connect qualitative and quantitative data, which can be done during data collection, analysis, or interpretation. For example, in a two-phase project, mining software repositories can be followed by a gender inference approach to support the selection of women’s data for a subsequent survey, connecting the two phases, as in (Qiu et al., 2019). In some cases, one form of data may support another, embedding a secondary form within a larger study. The researcher may weigh the collection of one type of data while using the other type to provide supplementary information.
@@
Regarding metrics for inclusion, SE research has been advancing on measuring the sense of belonging (Trinkenreich et al., 2023, 2024), which is the extent to which individuals feel like they belong or fit in a given environment (Hagerty and Patusky, 1995). Belongingness is a theoretical concept that is hard to observe directly, but it can be asked through different manifest variables (questions) and grouped on a latent construct. There are different instruments in the literature to measure a sense of belonging. The instrument used to measure belongingness in Open Source Software, for example, was based on the concept of a sense of virtual community (Blanchard, 2007) (a community that mainly interacts online) and included questions about feelings of membership as a member of the team, being known by others and knowing who to ask for help, feeling valued and perceiving the team is like home (Trinkenreich et al., 2023).
@@
In addition to the methods considered, the actual implementation of these methods for research purposes can be improved. On the diversity lens, most of the literature related to minorities in SE and underrepresentation is still focused on gender (Trinkenreich et al., 2022b), race (Rodríguez-Pérez et al., 2021; Nadri et al., 2021), neurodiversity (Morris et al., 2015; Márquez et al., 2024). English confidence is a metric to evaluate inclusion for people who are non-native in English (Serebrenik et al., 2020) and can include multiple questions to include both written and spoken communication and both technical and social contexts (Trinkenreich et al., 2023). Socio-economic factors are also essential to be measured. For example, Goel et al. suggest that most research for end-user programming targets WEIRD (Western, Educated, Industrialized, Rich, and Democratic) users, while ignoring non-Weird populations that make up 85% of the world (Goel et al., 2023).
@@
We need more diversity aspects and intersection of those, which is going to be discussed in the next section.
The overarching research goal for this theme is as follows:
@@
Research Goal: Develop methodologies and metrics to effectively analyze diversity and inclusion in software engineering, making use of mixed methods and online community data.
@@
Benefits (Utopia) Harms (Dystopia) Research directions
Enhanced inclusion metrics Comprehensive understanding of community dynamics Tailored intervention strategies for diversity
The definition of inclusion not always clear Lack of comprehensive analysis tools Oversimplified interpretations of complex identities
Bridging qualitative and quantitative research Developing new theoretical frameworks and theories from the data Utilizing online community data for inclusiveness measures
@@ 4.2. Intersectionality
The concept of intersectionality describes the ways in which social categories of identity, difference, and disadvantage, e.g., gender, race, ethnicity, sexual orientation, gender identity, disability, class, age, and other forms of discrimination, “intersect” simultaneously to create unique dynamics and effects (Cole, 2009). Intersectionality suggests that different diversity aspects are not mutually exclusive and do not operate in isolation. Research also shows the negative consequences for individuals at the intersection of diversity categories in SE and computing. For example, Ross et al. show that fewer black women are introduced to CS than non-Black women and Black men (Ross et al., 2020). In this case, Black women do not know whether their negative experiences should be attributed to their gender or race. Similarly, a recent study (van Breukelen et al., 2023) shows that older women developers adopt various “survival strategies” to persist in the tech industry, and are uncertain whether their negative experiences in software development environments are due to ageism (Baltes et al., 2020) or sexism.
Black women and veteran women are merely two possible intersections to consider when studying SDDI from an intersectional perspective. Many individuals also find themselves at the intersection of more than two diversity axes. Research suggests that White, able-bodied, and heterosexual male STEM professionals experience favorable treatment, while people with more intersections face reduced social inclusion, professional respect, career opportunities, salaries, and persistent intentions (Cech, 2022). Therefore, more work is needed to understand the experiences of developers at the intersection of multiple diversity aspects.
@@
Benefits (Utopia) Harms (Dystopia) Research Directions
In-depth understanding of diverse identities Enhanced methodologies for capturing intersectionality Empowerment through tailored interventions
Potential for oversimplification in the analysis Danger of marginalization through one-size-fits-all policies and analysis Ethical concerns in data collection and analysis
Developing a two-stage research approach to utilize both qualitative and quantitative data Understanding underexplored intersections of diversity in SE Designing interventions that respect and enhance self-perception and self-efficacy Adapting to AI-powered development and research environments
@@ 4.3. Knowledge Transfer
Substantial research activity is already ongoing in SE and beyond, targeting the effect of various diversity dimensions on the workforce. For example, existing work shows that masculine cultures can alienate women developers (Faulkner, 2007; Girls Who Code, 2019)—yet, studies show increased inclusion of women in development teams can enhance productivity (Vasilescu et al., 2015), community (Catolino et al., 2019) and code quality (Terrell et al., 2017). However, the transfer of these findings to other actors is ultimately vital to enhancing software development and software quality. Table 3 summarizes the harms of ignoring knowledge transfer, benefits of successful knowledge transfer, and research directions for the coming years.
@@
Benefits (Utopia) Harms (Dystopia) Research Directions
Seamless integration of research into industry practices Enhanced innovation through collaborative efforts Better alignment of academic curriculum with industry needs
Fragmented and siloed knowledge pools Industry and educational practices disconnected from current research and vice versa in the area of SDDI Overlap in the actions, actions not visible to others
Establishing frameworks for continuous exchange between academia and industry Cultivating partnerships for mutual knowledge enhancement Recognizing key actors from all areas of the quadruple helix
@@
In the last decades, SE research has increasingly tried to show industrial relevance in published work. This is witnessed by an increasing amount of publications with joint academic and industrial authors, special forums for industry-relevant work, such as the SE in Practice track at the International Conference on Software Engineering (ICSE), [^6] https://conf.researchr.org/track/icse-2025/icse-2025-software-engineering-in-practice a special issue in IEEE Software on sustaining software engineering knowledge transfer, [^7] https://www.computer.org/digital-library/magazines/so/call-for-papers-special-issue-on-sustaining-software-engineering-knowledge-transfer or funding calls that require collaboration between academia and industry (Business Finland, [n. d.]; Australian Research Council, [n. d.]). This focus on industry-relevant research has led to substantial work on how to transfer technology and knowledge, typically from academia to industry (see, e.g., (Gorschek et al., 2006)). However, we explicitly question whether the transfer of SE research knowledge has indeed been successful.
@@
Considering SDDI topics, we further believe that focusing solely on academia and industry for knowledge transfer is insufficient. SDDI initiatives are widespread beyond academia and industry—and fragmented. For example, government bodies pass legislation related to diversity and representation, [^8] For instance, the European Accessibility Act and the US Americans with Disabilities Act (ADA), Section 508 and societal groups promote specific aspects of diversity [^9] For instance, ACM-W, which is ”supporting, celebrating, and advocating for Women in Computing”. and techniques toward a more inclusive society. Thus, knowledge transfer of SDDI research is essential for each of the different actors in the quadruple helix: society, academia, government, and industry (Carayannis and Campbell, 2009). As a result of this spread and fragmentation, encouraging actions and initiatives may overlap and may be invisible to other actors. Therefore, we argue that research that considers SDDI in SE needs to engage with all four areas in the quadruple helix.
@@
Finally, re-considering the direction of knowledge transfer is important for research related to SDDI topics. In addition to transferring research results from academia to industry, society, and governments, researchers need to improve their knowledge of initiatives and the results obtained in the broader societal context. This also relates to the Methodology and Metrics theme, as methods that include stakeholders could be beneficial to reach this goal, e.g., participatory research or co-creation.
Overall, we summarize our discussion on this theme in the following research goal:

### v2

@@ 4. Results: Avoiding Undesirable Scenarios
In Section 3, we have outlined four scenarios on how SE industrial practice and higher education could look in 2030. These scenarios do not have a direct connection to SE research. However, SE research can and should contribute to industrial practice and higher education in order to avoid the undesirable scenarios and instead achieve a future closer to the desirable scenarios we described. To do so, we outline four orthogonal areas in which we believe SE research needs to evolve in the coming years.
@@
In Scenario S1, several SDDI metrics guided the company’s interventions toward a safer place for Jamie to develop their career. Similarly, Kris used various indicators in their educational setting in S2. If the university in Scenario S4 had monitored their students’ experience and well-being using suitable metrics, they would have noticed the lack of belonging and feeling of inclusion among students such as Moss. These three scenarios motivate the use of diversity metrics that can be used by practitioners and educators to better understand diversity in their environment and make decisions based on them. These metrics could range from obtaining demographic data, such as gender, country or region of birth, or age, to using various instruments to consider aspects such as personal preferences and interests, or cognitive diversity. Gender, in particular, has been one of the most studied aspects of diversity in software engineering, but the range of diversity aspects in the field has expanded recently (Rodríguez-Pérez et al., 2021). SE research can contribute to this area by developing appropriate Methodologies and Metrics.
@@
In Scenario S3, Ash’s experiences conflated frustrations due to multiple diversity dimensions: socioeconomic background, language background, and gender. In contrast, various intersectional backgrounds were effectively supported in Scenario S2 through personalized learning, which has been shown to reduce educational inequality and foster inclusion for learners from diverse cultures (Strekalova-Hughes et al., 2021). These scenarios highlight the importance of developing a better understanding of Intersectionality in SE research (Cech, 2022), that is, when multiple diversity aspects overlap.
@@
Both Scenarios S1 and S3 demonstrate the importance of transferring knowledge about SDDI to and from the SE field. For example, in Scenario S3, a lack of transfer of new ideas from outside means that traditional gender roles dominate the company culture, impacting Ash’s sense of belonging as they do not identify with these traditional gender roles and as their career progression is affected. Research has shown that women are still underrepresented in the industry, they still leave the industry as they do not feel a sense of belonging or are mistreated, or they even enter the field later in life as they lack guidance and support in choosing IT careers at a younger age. (Oliveira et al., 2024; McKinsey & Company, 2023b; Hyrynsalmi et al., 2024). In Scenario S1, Jamie and their employer thrive due to the adoption of diversity initiatives. This aligns with the findings that a diversity perspective is often missing in innovation and entrepreneurship policies, which predominantly target male-dominated industries. As such, there needs to be an increased push towards Knowledge Transfer between SE research, industry, education, government bodies, and societal groups to ensure diversity initiatives are effectively integrated into innovation policies and practices. (Liberda and Zajkowska, 2017)
@@
Finally, industrial practice and higher education do not operate in isolation but are embedded in the larger socioeconomic context. For example, in Scenarios S1 and S3, factors like income or government funding had a decisive effect on the workplace, highlighting differences in how investments shape outcomes in high- and low-income regions, as research has shown (Okoye et al., 2022). Therefore, it is imperative to use better existing Large-scale Socioeconomic Understandings developed by other disciplines in SE research to provide more inclusive workplace environments and higher education environments.
@@ 4.1. Methodologies and Metrics
Methodologies and metrics are essential for scientific exploration and discovery (Creswell and Creswell, 2017). Methodologies entail systematic research approaches for acquiring new knowledge through various research methods for data collection and analysis (Wohlin and Runeson, 2021). Metrics are used to quantify a characteristic of a real-world entity (Basili, 1988). SE research, in particular, uses a variety of methods and metrics to provide empirical evidence to support software development processes and tools (Easterbrook et al., 2008). For example, previous work has used different techniques and measures to explore concepts related to SDDI, including surveys to understand developers’ sense of belonging (Trinkenreich et al., 2023, 2024), interviews to understand barriers for women in online programming communities (Ford et al., 2016), and studies mining GitHub repositories to explore toxicity in open source software (Miller et al., 2022). However, while all these examples are related to SDDI, an underlying theory describing how they connect for SDDI and what is the consequence for the software being produced is missing. For example, how do we investigate toxicity as a barrier to minorities’ participation and belonging to a software development project? “How” includes methods and metrics.
The Linux Foundation’s Community Health Analytics in Open Source Software (CHAOSS) [^9] https://chaoss.community/unveiling-the-impact-dei-metrics-overcoming-social-barriers-in-open-source/ has identified several key metrics to enhance diversity. For example: (1) metrics for diversifying recruitment which focus on community engagement and labor investment, promoting equal opportunities and assessing candidates’ value; (2) metrics for leadership, which serve as key performance indicators, enhancing diversity and guiding strategies for inclusive participation while preventing authoritarian dynamics; (3) metrics on psychological safety and burnout which create environments where participants can raise their concerns and receive support; (4) mentorship and retention metrics which encourage collaboration and track new member engagement.
Table 1 summarizes potential benefits, harms and research directions for this theme.
@@
The focus of this theme is not only on what has been published in the field of SE around diversity and inclusion and with which methods—but also a look at multiple domains, that is, SE, educational psychology, and management, to investigate more how insights from these fields could be incorporated into the methodologies and metrics used to research SDDI in SE.
4.1.1. Methods
SDDI research often employs mixed methods (Creswell and Creswell, 2017), utilizing qualitative and quantitative data. Using this data more efficiently is an important future goal. Weighting, timing, and mixing are aspects that must be considered before planning mixed-methods SDDI research.
Weighting refers to the priority given to qualitative and quantitative data in research (Creswell and Creswell, 2017). The priority can be the same or favor one over another, depending on the researcher’s goals and the audience. For instance, if the aim is to advocate for policy changes, quantitative metrics may take precedence to provide compelling evidence to decision-makers. Conversely, if the focus is on fostering a supportive community, qualitative insights might be prioritized to understand personal experiences better.
@@ 4.1.1. Methods
Timing is about deciding whether to collect data sequentially in phases or concurrently (Creswell and Creswell, 2017). In sequential data collection, either qualitative or quantitative data collection can occur first, depending on the research goal. Concurrent data collection involves simultaneously gathering data for qualitative and quantitative analyses, which can be interesting in time-sensitive projects where contacting participants multiple times for data collection is not feasible (e.g., single surveys with open and closed-ended questions) (Trinkenreich et al., 2022a). Previous SDDI mixed-methods research used both concurrent and sequential mixed-methods research. Trinkenreich et al. (2022a) followed a concurrent mixed-methods research collecting data through a single survey to qualitatively uncover the challenges faced by women in software development teams and segment those challenges across demographics of age, caregiving responsibilities, marital status, and tenure. Examples of SDDI sequential mixed-methods research included surveys and mining software repository studies in varying orders. Vasilescu et al. (2015) started with a survey on perceptions of team diversity and then mined a software repository to measure how team productivity and turnover are impacted by gender and tenure diversity. Following the opposite order, Prana et al. (2021) started mining software repositories to investigate differences in gender inclusion in projects across geographical regions, followed by a survey aimed at developers from the various regions about factors that can potentially contribute to differences in developer participation based on gender and geography worldwide.
@@
Mixing involves choosing how to integrate or connect qualitative and quantitative data, which can be done during data collection, analysis, or interpretation. For example, in a two-phase project, mining software repositories can be followed by a gender inference approach to support the selection of women’s data for a subsequent survey, connecting the two phases, as done by Qiu et al. (Qiu et al., 2019). In some cases, one form of data may support another, embedding a secondary form within a larger study. For example, researchers might primarily collect quantitative data through surveys measuring employees’ stress levels and job satisfaction. To enhance these findings, they could embed qualitative interviews where employees share personal stories and experiences related to their well-being. The researcher may weigh the collection of one type of data while using the other type to provide supplementary information.
4.1.2. Metrics
Measuring diversity- and inclusion-related phenomena is a challenging task due to the need to capture complex social constructs, e.g., gender identity, sense of belonging or suitability of technology, being mindful that these constructs are highly likely to be perceived and conceptualized very differently by marginalized populations and recognizing the need to focus on the experiences of these populations (Ford and Johnson, 2024).
At the level of demographics, gender, and in particular opposition between women and men, is the most studied diversity aspect in software engineering (Rodríguez-Pérez et al., 2021). However, despite the extensive usage of this construct, acquiring information about the gender identity of study participants—whether in controlled experiments, interviews, and surveys—or individuals contributing to datasets used in data-driven archival studies, such as repository mining, presents significant challenges. Inadequate methods for gathering this information risk both alienating participants and undermining the validity of the scientific findings (Serebrenik, 2024). Even more so, given the well-known gender-related differences in information processing (Burnett et al., 2016b), one should ask themselves whether gender is an appropriate construct to study as opposed to one of the information processing facets, that can be reliably assessed by means of a validated instrument (Hamid et al., 2024).
@@ 4.1.2. Metrics
Finally, regarding metrics for inclusion, SE research has been advancing on measuring the sense of belonging (Trinkenreich et al., 2023, 2024), which is the extent to which individuals feel like they belong or fit in a given environment (Hagerty and Patusky, 1995). Belongingness is a theoretical concept that is hard to observe directly, but it can be asked through different manifest variables (questions) and grouped on a latent construct. There are different instruments in the literature to measure a sense of belonging. The instrument used to measure belongingness in Open Source Software, for example, was based on the concept of a sense of virtual community (Blanchard, 2007) (a community that mainly interacts online) and included questions about feelings of membership as a member of the team, being known by others and knowing who to ask for help, feeling valued and perceiving the team is like home (Trinkenreich et al., 2023).
@@
In addition to the methods considered, the actual implementation of these methods for research purposes can be improved. On the diversity lens, most of the literature related to minorities in SE and underrepresentation is still focused on gender (Canedo et al., 2020; Trinkenreich et al., 2022b; van Breukelen et al., 2023), race (Rodríguez-Pérez et al., 2021; Nadri et al., 2021), neurodiversity (Morris et al., 2015; Márquez et al., 2024). English confidence is a metric to evaluate inclusion for people who are non-native in English (Serebrenik et al., 2020) and can include multiple questions to include both written and spoken communication, and both technical and social contexts (Trinkenreich et al., 2023). Socio-economic factors are also essential to be measured. For example, Goel et al. suggest that most research for end-user programming targets WEIRD (Western, Educated, Industrialized, Rich, and Democratic) users, while ignoring non-WEIRD populations that make up 85% of the world (Goel et al., 2023).
@@
We need more diversity aspects and intersection of those, which is going to be discussed in the next section. The overarching research goal for this theme is as follows:
@@
Research Goal: Develop methodologies and metrics to effectively analyze diversity and inclusion in software engineering, making use of mixed methods and community data.
@@
Benefits Harms Research Directions
Enhanced inclusion metrics Comprehensive understanding of community dynamics Tailored intervention strategies for diversity The definition of inclusion not always clear Lack of comprehensive analysis tools Oversimplified interpretations of complex identities Bridging qualitative and quantitative research Developing new theoretical frameworks and theories from the data Utilizing online community data for inclusiveness measures
@@ 4.2. Intersectionality
Introduced by Kimberlé Crenshaw (Crenshaw, 1988), the concept of intersectionality describes the ways in which social categories of identity, difference, and disadvantage, e.g., gender, race, ethnicity, sexual orientation, gender identity, disability, class, age, and other forms of discrimination, “intersect” simultaneously to create unique dynamics and effects (Cole, 2009). Intersectionality suggests that different diversity aspects are not mutually exclusive and do not operate in isolation. Research also shows the negative consequences for individuals at the intersection of diversity categories in SE and computing. For example, Ross et al. show that fewer black women are introduced to CS than non-Black women and Black men (Ross et al., 2020). In this case, Black women do not know whether their negative experiences should be attributed to their gender or race. Similarly, a recent study by van Breukelen et al. (van Breukelen et al., 2023) shows that older women developers adopt various “survival strategies” to persist in the tech industry, and are uncertain whether their negative experiences in software development environments are due to ageism (Baltes et al., 2020) or sexism.
Black women and older women are merely two possible intersections to consider when studying SDDI from an intersectional perspective. Many individuals also find themselves at the intersection of more than two diversity axes. Research suggests that White, able-bodied, and heterosexual male STEM professionals experience favourable treatment, while people with more intersections face reduced social inclusion, professional respect, career opportunities, salaries, and persistent intentions (Cech, 2022). Therefore, more work is needed to understand the experiences of developers at the intersection of multiple diversity aspects. While this need has been recognized, the number of such studies remains limited (Ford et al., 2019; Sánchez-Gordón and Colomo-Palacios, 2024; Kohl and Prikladnicki, 2024; Szlávi et al., 2023).
@@
Benefits Harms Research Directions
In-depth understanding of diverse identities Enhanced methodologies for capturing intersectionality Empowerment through tailored interventions Potential for oversimplification in the analysis Danger of marginalization through one-size-fits-all policies and analysis Ethical concerns in data collection and analysis Developing a two-stage research approach to utilize both qualitative and quantitative data Understanding underexplored intersections of diversity in SE Designing interventions that respect and enhance self-perception and self-efficacy Adapting to AI-powered development and research environments
@@ 4.3. Knowledge Transfer
Substantial research activity is already ongoing in SE and beyond, targeting the effect of various diversity dimensions on the workforce. For example, existing work shows that masculine cultures can alienate women developers (Faulkner, 2007; Girls Who Code, 2019)—yet, studies show increased inclusion of women in development teams can enhance productivity (Vasilescu et al., 2015), community (Catolino et al., 2019) and code quality (Terrell et al., 2017). However, transferring these findings to other actors is ultimately vital to enhancing software development and quality. Overall, co-creation or transferring findings and best practices between different actors such as industry, academia, government, and society is becoming more crucial, as technology has more cross-cutting impact in every sector. Therefore, we approach the topic of knowledge transfer in advancing SDDI. Table 3 summarizes the benefits of successful knowledge transfer, the harms of ignoring knowledge transfer, and research directions for the coming years.
@@
Benefits Harms Research Directions
Seamless integration of research into industry practices Enhanced innovation through collaborative efforts Better alignment of academic curriculum with industry needs Fragmented and siloed knowledge pools Industry and educational practices disconnected from current research and vice versa in the area of SDDI Overlap in the actions, actions not visible to others Establishing frameworks for continuous exchange between academia and industry Cultivating partnerships for mutual knowledge enhancement Recognizing key actors from all areas of the quadruple helix
@@
In the last decades, SE research has increasingly paid attention to showing industrial relevance in published work. This is witnessed by an increasing amount of publications with joint academic and industrial authors, special forums for industry-relevant work, such as the SE in Practice track at the International Conference on Software Engineering (ICSE), [^10] https://conf.researchr.org/track/icse-2025/icse-2025-software-engineering-in-practice a special issue in IEEE Software on sustaining software engineering knowledge transfer, [^11] https://www.computer.org/digital-library/magazines/so/call-for-papers-special-issue-on-sustaining-software-engineering-knowledge-transfer or funding calls that require collaboration between academia and industry (Business Finland, [n. d.]; Australian Research Council, [n. d.]). This focus on industry-relevant research has led to substantial work on transferring technology and knowledge, typically from academia to industry (see, e.g., (Gorschek et al., 2006)). However, we explicitly question whether the transfer of SE research knowledge, especially, has indeed been successful.
@@
Considering SDDI topics, we further believe that focusing solely on academia and industry for knowledge transfer is insufficient. SDDI initiatives are widespread beyond academia and industry—and fragmented. For example, government bodies pass legislation related to diversity and representation [^12] For instance, the European Accessibility Act and the US Americans with Disabilities Act (ADA), Section 508, and societal groups promote specific aspects of diversity [^13] For instance, ACM-W, which is ”supporting, celebrating, and advocating for Women in Computing”. and techniques toward a more inclusive society. Thus, knowledge transfer of SDDI research is essential for each of the different actors in the quadruple helix (Figure 3): society, academia, government, and industry (Carayannis and Campbell, 2009). We have chosen the quadruple helix approach to visualize all the actors in our field: In the Quadruple Helix model, society refers to citizens, media, cultural organizations, and non-profit organizations. In the software engineering field, these kinds of actors could be media outlets such as TechCrunch and Wired or non-profit communities and organizations such as ’Girls Who Code’ or Czechitas (Buhnova, 2024). In business, they refer to software engineering companies, and in government, they refer to the public policy around technology or public funding for software companies and non-profit organizations. In academia, it refers to universities, research centres and conferences.
As a result of this spread and fragmentation, encouraging actions and initiatives may overlap and remain invisible to other actors. More efficient knowledge transfer between different actors could benefit all, particularly as some technology companies are reducing their DEI teams  (Burnett and Aguinis, 2024; Alfonseca and Zahn, 2023). By engaging actors from academia, industry, government, and society, we could promote more efficient inclusive mentorship practices that foster the inclusion of women and minorities in professional and OSS contexts (Jacobs et al., 2024; Hyrynsalmi and Sutinen, 2019). Up-to-date knowledge from different sectors can also prevent situations in which well-meant policies can lead to unintended negative consequences or tokenism (Burnett and Aguinis, 2024).
Figure 3. Quadruple helix
Therefore, we argue that research that considers SDDI in SE needs to engage with all four areas in the quadruple helix. To aim for an even more ambitious knowledge transfer level, we might even consider including the ’environment’ in knowledge transfer, i.e., the sustainability aspect. This leads to a ’Quintuple helix’, which is especially focused on enabling sustainable and environmentally conscious innovation, as the quadruple helix emphasizes innovation driven by collaboration between the private sector, government, academia, and civil society (Carayannis and Campbell, 2009; Carayannis et al., 2012).
@@
Finally, re-considering the direction of knowledge transfer is important for research related to SDDI topics. In addition to transferring research results from academia to industry, society, and governments, researchers need to improve their knowledge of initiatives and the results obtained in the broader societal context. This also relates to the Methodology and Metrics theme, as methods that include stakeholders could be beneficial to reach this goal, e.g., participatory research or co-creation. Overall, we summarize our discussion on this theme in the following research goal: